40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Feb 13 16:35:32 2020, anchal, DailyProgress, BEAT, BN Detector was saturated. Reduced laser powers. 
    Reply  Fri Feb 14 10:59:13 2020, anchal, DailyProgress, BEAT, Beatnote noise lowered in low frequency region 
       Reply  Mon Feb 17 16:08:23 2020, anchal, DailyProgress, BEAT, Beatnote noise lowered in low frequency region with ISS CTN_Noise_Budget_Effect_ISS_ON.pdf
          Reply  Fri Feb 21 16:22:04 2020, anchal, DailyProgress, BEAT, Beatnote noise lowered in low frequency region with ISS CTN_Noise_Budget_Effect_ISS_ON_New.pdf
             Reply  Tue Mar 3 10:02:12 2020, anchal, DailyProgress, BEAT, Lowest ever beatnote spectrum today! CTN_Noise_Budget_March03-2020.pdf
                Reply  Thu Mar 12 11:06:35 2020, anchal, DailyProgress, BEAT, Lowest ever beatnote spectrum today! CTN_Noise_Budget_March11-2020.pdf
                   Reply  Wed Mar 18 12:51:22 2020, anchal, DailyProgress, BEAT, Super beatnote measurement analysis results CTN_Beatnote_SuperMeasurement_March13-16.pdf
                      Reply  Mon Apr 13 16:52:30 2020, anchal, DailyProgress, BEAT, Beatnote measurements back on track CTN_Latest_BN_Spec.pdf
Message ID: 2566     Entry time: Mon Apr 13 16:52:30 2020     In reply to: 2562
Author: anchal 
Type: DailyProgress 
Category: BEAT 
Subject: Beatnote measurements back on track 

Since this morning atleast, I'm not seeing the North Path unstability (see CTN:2565) and the beatnote is stable and calm at the setpoint. Maybe the experiment just needed some distance from me for few days.

So today, I took a general single shot measurement and even after HEPA filers being on at 'Low', the measurement is the lowest ever, especially in the low-frequency region. This might be due to reduced siesmic activity around the campus. I have now started another super beatnote measurement which would take measurement continuously every 10 min is the transmission power from the cavities look stable enough to the code.

there is a new broad bump though arounf 250-300 Hz which was not present before. But I can't really do noise hunting now, so will just take data until I can go to the experiment.


Latest BN Spectrum: CTN_Latest_BN_Spec.pdf

Daily BN Spectrum: CTN_Daily_BN_Spec.pdf


Relevant post:

CTN:2565: North path's buggy nature NOT solved

Attachment 1: CTN_Latest_BN_Spec.pdf  164 kB  Uploaded Mon Apr 13 17:58:11 2020  | Hide | Hide all
CTN_Latest_BN_Spec.pdf
ELOG V3.1.3-